$SPAD/src/input solveperf.input

Raymond E. Rogers December 3, 2016

Abstract

The last call to solve is commented out. It takes about 10 minutes to run. I do not know why.

Contents

2

— * —

)set break resume
)sys rm -f solveperf.output
)spool solveperf.output
)set message test on
)set message auto off
)clear all
)set mes time on
--S 1 of 9
list:=[p,Vr,Vt,e]
--R
--R
--R   (1)  [p,Vr,Vt,e]
--R                                 Type: List(OrderedVariableList([p,Vr,Vt,e]))
--I                                                   Time: 0.04 (OT) = 0.04 sec
--E 1
--S 2 of 9
eq1a:=((-Vr^3+Vr^2)*Vt+Vr^3-Vr^2)
--R
--R
--R             3     2        3     2
--R   (2)  (- Vr  + Vr )Vt + Vr  - Vr
--R                                                    Type: Polynomial(Integer)
--I                           Time: 0.02 (EV) + 0.04 (OT) + 0.02 (GC) = 0.08 sec
--E 2
--S 3 of 9
eq1:=((-Vr^3+Vr^2)*Vt+Vr^3-Vr^2)*R*p
--R
--R
--R                3       2          3       2
--R   (3)  ((- R Vr  + R Vr )Vt + R Vr  - R Vr )p
--R                                                    Type: Polynomial(Integer)
--I                                                                  Time: 0 sec
--E 3
--S 4 of 9
eq2:=Vr*Vt^3+Vr*Vt^2-Vr*Vt+Vr*p
--R
--R
--R                    3        2
--R   (4)  Vr p + Vr Vt  + Vr Vt  - Vr Vt
--R                                                    Type: Polynomial(Integer)
--I                                                   Time: 0.02 (IN) = 0.02 sec
--E 4

--S 5 of 9
eq3:=(R*Vr*Vt+R*Vr)*e+((R-1)*Vr+1)*Vt-R*Vr
--R
--R
--R   (5)  (R Vr Vt + R Vr)e + ((R - 1)Vr + 1)Vt - R Vr
--R                                                    Type: Polynomial(Integer)
--I                                                   Time: 0.01 (OT) = 0.01 sec
--E 5
--S 6 of 9
eq4:=-r^2+(Vt^2)+(Vr^2)
--R
--R
--R           2     2     2
--R   (6)  - r  + Vt  + Vr
--R                                                    Type: Polynomial(Integer)
--I                                                                  Time: 0 sec
--E 6
--S 7 of 9
Eqlista:=[eq1a,eq2,eq3,eq4]
--R
--R
--R   (7)
--R         3     2        3     2              3        2
--R   [(- Vr  + Vr )Vt + Vr  - Vr , Vr p + Vr Vt  + Vr Vt  - Vr Vt,
--R                                                     2     2     2
--R    (R Vr Vt + R Vr)e + ((R - 1)Vr + 1)Vt - R Vr, - r  + Vt  + Vr ]
--R                                              Type: List(Polynomial(Integer))
--I                                                                  Time: 0 sec
--E 7
--S 8 of 9
Eqlist:=[eq1,eq2,eq3,eq4]
--R
--R
--R   (8)
--R            3       2          3       2                3        2
--R   [((- R Vr  + R Vr )Vt + R Vr  - R Vr )p, Vr p + Vr Vt  + Vr Vt  - Vr Vt,
--R                                                     2     2     2
--R    (R Vr Vt + R Vr)e + ((R - 1)Vr + 1)Vt - R Vr, - r  + Vt  + Vr ]
--R                                              Type: List(Polynomial(Integer))
--I                                                                  Time: 0 sec
--E 8
--S 9 of 9
solve(Eqlista,list)
--R
--R

--R   (9)
--R   [
--R                   4       2                              2
--R         (- e - 1)r  + 4e r  - 4e + 2             (e + 1)r  - 2e
--R     [p= ----------------------------, Vr= 1, Vt= --------------,
--R                       2                                 2
--R        2           2     2
--R      (e  + 2e + 1)r  - 2e  - 2= 0]
--R     ,
--R
--R                               2
--R     [p= - 1, Vr= (- 2R e + 1)r  + 2R e - 1, Vt= 1,
--R         2 2             2     2 2
--R      (4R e  - 4R e + 1)r  - 4R e  + 4R e - 2= 0]
--R     ]
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--I               Time: 0.02 (IN) + 0.87 (EV) + 0.07 (OT) + 0.11 (GC) = 1.07 sec
--E 9
-- This takes a long time. Figure out why.
--
--solve(Eqlist,list)
--

)spool
)lisp (bye)

———-

References

[1] http://axiom-wiki.newsynthesis.org/SandBoxShortcoming